OpenStudy (anonymous):

ok i did what you said for this question Choose the correct slope of the line that passes through the points (−6, −6) and (−3, 1). nd i got -5/-9 what did i do wrong ?

OpenStudy (anonymous):

slope = \[(y2-y1)/(x2-x1)\] so, slope = ((1)-(-6))/((-3)-(-6)) and therefore, slope=7/3